OpenStudy (anonymous):

slope 2; y-intercept (0, -3)

OpenStudy (anonymous):

need help, please

OpenStudy (anonymous):

1) Write general equation for lines: y = mx + n (m is slope) 2) Put slope = 2 y = 2x + n 3) Put the values of the point you have (x,y) = (0, -3) -3 = 2*0 + n n = -5 4) Write first equation and put values of m and n into the equation y = 2x -5

OpenStudy (anonymous):

how does n=-5

OpenStudy (anonymous):

Sorry :D n = -3 y = 2x-3